### v1

# Generative Modeling of Neural Dynamics via Latent Stochastic Differential Equations

Abstract
We propose a probabilistic framework for developing computational models of biological neural systems. In this framework, physiological recordings are viewed as discrete-time partial observations of an underlying continuous-time stochastic dynamical system which implements computations through its state evolution. To model this dynamical system, we employ a system of coupled stochastic differential equations with differentiable drift and diffusion functions and use variational inference to infer its states and parameters. This formulation enables seamless integration of existing mathematical models in the literature, neural networks, or a hybrid of both to learn and compare different models. We demonstrate this in our framework by developing a generative model that combines coupled oscillators with neural networks to capture latent population dynamics from single-cell recordings. Evaluation across three neuroscience datasets spanning different species, brain regions, and behavioral tasks show that these hybrid models achieve competitive performance in predicting stimulus-evoked neural and behavioral responses compared to sophisticated black-box approaches while requiring an order of magnitude fewer parameters, providing uncertainty estimates, and offering a natural language for interpretation.
1 Introduction
Unlike many scientific fields, neuroscience lacks a unified theoretical framework that bridges micro-scale mechanisms with macro-scale observations. That is, while the biophysical processes underlying individual neuron spiking have been well-understood for decades (Hodgkin & Huxley, 1952), the computational principles governing collective neural activity that give rise to cognition and behavior remain elusive.
Dynamical systems theory provides a promising framework to address this gap (Durstewitz et al., 2023). Within this formalism, the brain is viewed as a complex dynamical system where neural activity patterns evolve through a high-dimensional state space according to well-defined governing equations, providing a mathematical language for understanding how distributed neural circuits implement computation and generate behavior (Van Gelder, 1998; Izhikevich, 2007; Deco et al., 2008; Breakspear, 2017; Favela, 2021). This perspective has led to diverse mathematical models, ranging from mean-field approximations of biophysical models that preserve mechanistic interpretability (Wilson & Cowan, 1972; Jirsa & Haken, 1996; Robinson et al., 1997), to phenomenological models inspired by statistical physics that capture emergent collective behavior (Amari, 1977; Hopfield, 1982; Sompolinsky et al., 1988). However, while these models have provided key insights into various neural phenomena (Goldman-Rakic, 1995; Wang, 2002; Fries, 2005), fitting them to single-trial neural data has proven challenging (Urai et al., 2022).
In recent years, latent variable models have emerged as a powerful alternative. These models represent high-dimensional neural activity in terms of low-dimensional latent states evolving according to Markovian dynamics (Paninski et al., 2010; Hurwitz et al., 2021a). Neural computations are thus assumed to be implemented through dynamic motifs in a low-dimensional latent space (e.g., attractors, oscillations, bifurcations) (Vyas et al., 2020; Khona & Fiete, 2022; Durstewitz et al., 2023). The popularity of this approach stems from converging empirical evidence that neural population activity during simple tasks lives on a low-dimensional manifold (Gao & Ganguli, 2015). Coupled with ongoing developments in machine learning (Kingma & Welling, 2013; Fabius & Van Amersfoort, 2014; Rubanova et al., 2019), latent variable models are becoming a standard tool in systems and computational neuroscience (Chen et al., 2018; Glaser et al., 2020; Hurwitz et al., 2021b; Zhou & Wei, 2020; Kim et al., 2021).
However, selecting an appropriate model for representing latent neural dynamics presents a significant challenge. The model must be expressive enough to capture complex non-linear dynamics, supporting flexible neural computations, while remaining amenable to interpretation and analysis to be useful for hypothesis generation and testing. Furthermore, model predictions should account for various sources of uncertainty typically present in such modeling endeavors, such as measurement noise, process noise, and model uncertainty. Finally, to utilize the inferred model for online applications as in brain computer interfaces (BCIs), there should be an efficient tractable method for sampling from the model in real-time.
We posit that recent developments in scientific machine learning, where domain-specific models — often expressed as differential equations — are combined with neural networks and scalable optimization techniques provide an untapped potential to address these challenges (Raissi et al., 2019; Rackauckas et al., 2020; Lai et al., 2021; Karniadakis et al., 2021). Additionally, ongoing innovations in training (neural) stochastic differential equations (SDEs) (Tzen & Raginsky, 2019a; Li et al., 2020; Zeng et al., 2023; Course & Nair, 2024), provide an opportunity to develop scalable probabilistic models in neuroscience, a field wherein the systems in question are inherently stochastic (Laing & Lord, 2009; Rolls & Deco, 2010) and the existing measurement tools only provide a coarse proxy of the system under study (Stevenson & Kording, 2011; Urai et al., 2022).
Contributions
We develop a general framework for modeling neural dynamics using latent SDEs with differentiable state- and input-dependent drift and diffusion functions. Within this framework, we introduce a specific generative model, combining coupled oscillators with neural networks, to capture neural population dynamics. We evaluate our approach across multiple experimental scenarios, demonstrating its effectiveness and efficiency in predicting stimulus-evoked neural and behavioral responses in both reaching and decision-making tasks. Through systematic comparison of latent ODE- and SDE-based models under varying noise levels and state dimensionality in simulated systems, we highlight the critical importance of modeling process noise; a factor often overlooked in current latent variable approaches. Finally, our results suggest that data-driven coupled oscillators could provide an expressive yet interpretable framework for understanding neural dynamics, while maintaining computational tractability.
2 Framework
Notations
Let denote a continuous path of latent neural states, which we observe through neural recordings (e.g., spike counts, fMRI time-series) denoted by , where each is observed at time . Additionally, we may observe behavioral responses (e.g., movement trajectories, choice responses) represented as , where each . In some settings, these observations are driven by external stimuli or task variables (e.g., visual stimuli, auditory cues) denoted by , where each . A single trial is thus defined as , where , and/or could be missing. Given a dataset consisting of trials, our objectives are to: (1) infer the underlying neural dynamics that generated the observations, represented by the posterior distribution , (2) learn a generative model capable of predicting stimulus-evoked responses, represented by the joint distribution , and (3) extract interpretable dynamical features that provide insights into neural computations.
2.1 Generative model
We model latent neural dynamics as a continuous-time stochastic process, specified as the solution to the following Itô SDE:
with , where is the probability distribution of the initial condition , and is an interpolated encoded representation of . Here, are basis functions chosen for the desired interpolation scheme and is an input encoder whose form depends on the input modality. The functions and are Lipschitz-continuous differentiable functions defining the drift and diffusion coefficients of the SDE, respectively. Finally, denotes a standard -dimensional Wiener process.
This SDE induces a path measure on the space of continuous paths which implicitly define the distribution . Samples from this distribution are obtained by solving the SDE with different realizations of and :
The continuous-time latent states are mapped to discrete-time observations through:
where and denote the observation models (e.g., Poisson model for spike counts, Gaussian model for continuous measurements), with and being differentiable functions mapping the latent state to the parameters of these distributions.
2.2 Inference and learning
The true parameters of our generative model are unknown and need to be learned from the data . Moreover, the exact posterior distribution is intractable to compute. Variational inference provides an efficient and scalable approach to tackle this problem. However unlike standard latent variable models, the latent state in our model are realizations of a continuous-time stochastic process. Thus we need to define a parametric distribution over paths rather than random-valued vectors.
Following (Tzen & Raginsky, 2019a, b; Li et al., 2020), we define the approximate posterior distribution via an augmented neural SDE. Specifically, here we define the approximate posterior path as realizations of the following SDE (we omit time indices for ease of notation):
where and is a continuous path encoding the observations, with being a modality-specific observation encoder. We define the distribution of the initial condition as , where and are neural networks, and . We further set to prevent over-parameterizing the initial conditions and rendering the dynamics redundant (Zabihi et al., 2019).
Here is a neural network representing the drift of the augmented SDE. Note that while the drift term is different from that of the generative model, the diffusion term is shared. As illustrated in (Tzen & Raginsky, 2019a), this specification enables leveraging Girsanov’s theorem (Girsanov, 1960) to obtain a tractable KL divergence between the the generative (prior) and augmented (approximate posterior) distributions:
where with and denoting the measures induced by the generative and augmented SDE, respectively.
Using this we can define an evidence lower bound (ELBO) on the conditional marginal likelihood of the observations:
The ELBO consists of three main terms: (1) the expected log-likelihood of the observations under the approximate posterior, which includes both neural recordings and behavioral responses, (2) the KL divergence between the initial state distributions, and (3) the path-wise KL divergence between the approximate posterior and prior processes derived via Girsanov’s theorem (see Appendix B for detailed derivation of the loss).
The model parameters and variational parameters can be jointly optimized by maximizing this ELBO using stochastic gradient descent. Both the expectation over and the path-wise integral in the KL divergence term are computed using numerical integration schemes suitable for SDEs (see Appendix C for implementation details).
This formulation enables us to simultaneously learn the parameters of the generative model while performing approximate posterior inference over the latent neural trajectories. The learned generative model can then be used for various online and offline downstream tasks including probabilistic prediction of stimulus-evoked responses and extraction of interpretable dynamical features.
3 Hybrid modeling
So far we have described a general framework for modeling the latent dynamics of a neural system from empirical observations. This framework allows us to leverage almost arbitrary differentiable functions to model the drift and diffusion terms of an SDE. A natural choice would be to use a neural SDE to act as the generative model because of their universal approximation properties. However, this comes at the cost of efficiency and interpetability. An alternative option, is to leverage prior models proposed in the literature and use neural networks either to approximate their parameters or to augment their dynamics. Such hybrid models can improve sample-efficiency, generalization performance and, are easier to interpret in comparison to models that rely only on neural networks (Rackauckas et al., 2020; ElGazzar & van Gerven, 2024). In this section, we detail a specific instantiation of the framework which combines a well-understood dynamical system and neural networks to obtain an expressive, interpretable and parameter-efficient generative model of neural population dynamics.
3.1 Stochastic nonlinear coupled oscillators
A well-known class of models that are used to study several biological and physical systems are coupled oscillators. Coupled oscillators provide a rich yet simplified language for studying complex interactions in systems exhibiting emergent behavior such as synchronization, pattern formation and phase transitions (Winfree, 1980).
We focus on a specific instantiation of coupled oscillators (Matthews et al., 1991) described by the following complex-valued ODE:
where represents the position of oscillators in the complex plane, the bifurcation parameters, denotes the natural frequency, and represents all-to-all coupling strength. This formulation captures the essential dynamics of coupled limit-cycle oscillators near a supercritical Hopf bifurcation, where in the weak coupling regime, the term determines the limit cycle radius. Despite its simplicity, this system can exhibit diverse emergent dynamical regimes including frequency locking, amplitude death, and chaos through variations in its parameters and the initial conditions.
Here we propose to use this system as the drift of our generative SDE in our framework. Specifically, we use the position of the oscillator as the latent state in the generative SDE. Additionally, we consider the coupling strength as a time-varying input dependent function of the external input .
Incorporating these updates and writing the model in terms of the real and imaginary parts of the oscillator , we obtain the following system of SDEs:
where , , are neural networks representing the input dependent connectivity and the diffusion terms for the real and imaginary states respectively. The trainable parameters of the generative SDE are thus the parameters of the neural networks as well as and . We refer to this dynamics model as coupled oscillator SDE (CO-SDE).
3.2 Modeling multiple interacting populations
Often, we have access to neural activity recorded from neurons across multiple brain regions, simultaneously. Given access to recordings from brain regions, we can extend the generative model by factorizing the latent state where the number of oscillators can vary depending on the population size and complexity. The observed activity of a population is then
where is a vector of rate parameters which depends on a population-specific transformation that maps the latent state to the firing rates of the population.
4 Experiments
4.1 Simulated data
We first sought to evaluate the efficacy of our framework on modeling a partially observed neural system with different levels of process noise. To this end, we generated spiking data from the following system:
where describes the population average firing rate of three neural populations driven by a sinusoidal input with randomly sampled frequency and phase. and describe the recurrent connectivity and input interaction matrices, respectively. The time scale of the network dynamics is given by and describes the magnitude of process noise in the system. The system output is given by representing the observed spike counts of 50 neurons generated by sampling from an inhomogeneous Poisson process with rates computed by affine transformation of the population average activity via the output matrix . We fit our latent SDE model on the spiking data by maximizing the Poisson-log likelihood of the observations given predicted rate and evaluated performance using bits per spike (bps), a metric that quantifies how well the predicted firing rates explain the observed spike patterns (Pei et al., 2021).
For our experiments, we varied the noise level to test the robustness of the model. We used a latent neural SDE with varying sizes of latent dimension and compared against two baselines: a latent ODE (Rubanova et al., 2019) and LFADS (Pandarinath et al., 2018).
As demonstrated in Figure 2a, the latent SDE model can better explain the data variance when using fewer latent dimensions compared to a latent ODE model. Increasing the latent dimensions beyond the true state dimensions does not significantly improve the model fit unlike the ODE-based model which stagnates at a lower absolute value and higher latent dimension. Additionally, as shown in Figure 2b, the latent SDE model is more robust to process noise compared to latent ODE and LFADS models. These results suggest that latent SDEs can be more reliable to model real-world data when the true state dimension is unknown and process noise can emerge from unobserved interactions. In this setting, a probabilistic initial state with deterministic dynamics, as in the case of latent ODEs and LFADS, may not be sufficient.
| Datasets | ||||||
| Latent Dynamics | # Params ↓ | Perturbed Reach | Visual Decision | Delayed Reach | ||
| Neural | Behavior | Neural | Behavior | Neural | ||
| (bps) ↑ | () ↑ | (bps) ↑ | () ↑ | (bps) ↑ | ||
| RNN | 8320 | 0.175 ± 0.008 | 0.84 ± 0.02 | 0.14 ± 0.015 | 0.59 ± 0.03 | 0.291 ± 0.012 |
| GRU | 24960 | 0.187 ± 0.004 | 0.81 ± 0.015 | 0.16 ± 0.008 | 0.63 ± 0.02 | 0.305 ± 0.006 |
| LSTM | 33280 | 0.191 ± 0.003 | 0.85 ± 0.01 | 0.16 ± 0.006 | 0.66 ± 0.015 | 0.309 ± 0.004 |
| Neural ODE | 4880 | 0.221 ± 0.005 | 0.86 ± 0.012 | 0.15 ± 0.01 | 0.59 ± 0.025 | 0.316 ± 0.008 |
| Coupled Oscillator ODE | 465 | 0.221 ± 0.006 | 0.85 ± 0.015 | 0.21 ± 0.009 | 0.61 ± 0.02 | 0.309 ± 0.007 |
| Neural SDE | 4944 | 0.243 ± 0.007 | 0.86 ± 0.015 | 0.18 ± 0.012 | 0.70 ± 0.02 | 0.317 ± 0.009 |
| Coupled Oscillator SDE | 526 | 0.238 ± 0.008 | 0.86 ± 0.018 | 0.22 ± 0.011 | 0.67 ± 0.025 | 0.314 ± 0.01 |
4.2 Empirical data
4.2.1 Datasets & baselines
We evaluated our framework across three neuroscience datasets spanning different brain regions and behavioral tasks.
Perturbed reach task In this experiment, a monkey is presented with a target location and tasked with moving the cursor from the center to the target location by controlling a manipulandum (Chowdhury et al., 2020) (Figure 3a). On one portion of the trials, the monkey’s arm was perturbed before the reach by applying resisting forces on the manipulandum. Neural activity was recorded from Brodmann’s area 2 of the somatosensory cortex. Here, we trained a model to predict stimulus-evoked neural responses in area 2 as well as the hand position of the monkey. The input to the model are the target location, as well as the perturbing force and torques on the manipulandum.
Visual decision making In this task, a mouse is presented with images of different contrasts and the mouse is tasked with turning a wheel to bring the image with higher contrast to the center of the screen when a go cue is randomly presented (Steinmetz et al., 2019) (Figure 3d). Neural activity is recorded in multiple regions across the brain using Neuropixels probes. Here, we trained the model to predict stimulus-invoked neural responses across several brain regions simultaneously as well as predict the wheel velocity. The input to the model are the contrast levels, the timing of their presentation, and the timing of the go cue.
Delayed reach task The task is a delayed center-out reach task with barriers, resulting in a variety of straight and curved trajectories (Mante et al., 2013) (Figure 3g). Neural activity was recorded from the dorsal premotor (PmD) and primary motor (M1) cortices, and cursor, monkey gaze position, and monkey hand position and velocity are also provided. Here we train a model to predict neural responses during movement in both areas given the preparatory activity before the go cue. No stimulus information is provided to the model during training or prediction.
We compare different latent variable models across the three datasets. To enable a fair compariosn, we fix the encoder and decoder architectures across all models while varying the latent sequential model. That is, we compare the performance of RNN, GRU, LSTM, neural ODE, and coupled oscillator ODE models with that of our proposed coupled oscillator SDE and a neural SDE model. All the baseline models had a probabilistic initial state and were trained via variational inference. Note that this formulation enables generalizing the baselines to several previous models proposed in literature (Pandarinath et al., 2018; Hurwitz et al., 2021b; Kim et al., 2021; Versteeg et al., 2023) (See Appendix D for details on the model architecture).
4.2.2 Empirical results
Table 1 presents 5-fold cross-validation results across the three datasets, evaluating both neural activity prediction (in bits per spike) and behavioral performance (via the coefficient of determination, ). First, we observe that our stochastic models (Neural SDE and CO-SDE) on average consistently outperform their deterministic counterparts. This suggests that incorporating stochasticity into the latent dynamics is crucial for capturing the variability inherent in neural and behavioral data and that a stochastic initial state might not be sufficient to capture uncertainty in the dynamics.
Second, we observe that model complexity, as measured by parameter count, does not correlate with performance. The coupled oscillator models achieve competitive or superior performance while using only 465-526 parameters, compared to 33,280 parameters for the LSTM model. Note that while this number does not account for the number of parameters required for modeling the approximate posterior SDE (3528), this remains a significant improvement in parameter efficiency for the generative model. This suggests that incorporating appropriate inductive biases through the coupled oscillator architecture enables more efficient modeling of the underlying dynamics. This also suggests that standard deep learning models may be too over-parameterized to learn the underlying dynamical system during simple behavioral tasks.
Beyond quantitative results, oscillator-based models provide a direct way to interpret latent neural dynamics. We visualize the value of the inferred frequencies of the model after training as well as the value of the coupling strength in response to sample stimuli in Figure 3. In the perturbed reach task, where we have data from one somatosensory area, we observe that the connectivity strength increases over time (Figure 3c). This is also observed in the delayed reach task where we have recordings from two regions in the motor in the motor cortex (Figure 3k). In contrast, on the visual decision making task where we have recordings from multiple regions across the brain the coupling strength decreases over time (Figure 3f). This decrease is marked via two significant drops at two timepoints which mark the presentation of the stimulus and the go cue, respectively. Additionally, we observe that on average, the inferred natural frequency of the oscillators differ between brain regions. Notably, the average frequency of oscillators corresponding to M1 is higher than the average frequency of oscillators corresponding to PMd during the delayed reach task (Figure 3k), which may be linked to distinct functional roles related to action preparation and execution (Thura et al., 2022).
5 Discussion
This work introduces a probabilistic framework for modeling neural dynamics that combines the expressiveness of neural networks with the interpretability of mechanistic or phenomenological models. Our empirical results across simulated and real datasets demonstrate several key advantages of this approach.
First, explicitly modeling probabilistic dynamics through SDEs consistently improves model performance compared to deterministic alternatives, even when the latter incorporate uncertainty through probabilistic initial conditions. This suggests that neural variability stems not just from uncertain initial states but from inherently stochastic dynamics (Laing & Lord, 2009). The superior performance of SDE-based models in high-noise regimes as shown in the simulations results further supports this conclusion. Additional simulation results also show that using stochastic dynamics can improve model fit while relying on lower-dimensional latent states; an observation that is also found in stochastic low-rank RNNs (Pals et al., 2024).
Second, our hybrid coupled oscillator model achieves competitive or superior performance while requiring an order of magnitude fewer parameters than standard deep learning approaches. This parameter efficiency suggests that incorporating appropriate inductive biases through mechanistic components can effectively constrain the model space (Rackauckas et al., 2020; ElGazzar & van Gerven, 2024). The success of coupled oscillators in particular hints at their potential as a natural description language for neural population dynamics, consistent with existing theoretical work on neural oscillations (Buzsaki & Draguhn, 2004; Breakspear et al., 2010; Churchland et al., 2012; Bick et al., 2020). This also supports the potential of coupled oscillators as an expressive and efficient tool to model arbitrary time-series data, as seen in emerging machine learning research (Rusch & Mishra, 2020; Effenberger et al., 2022; Rusch & Rus, 2024).
With that said, there are a number of limitations to the proposed approach which merits consideration. First, the framework relies on variational inference for training the model parameters. This could be suboptimal as variational inference is known to provide overconfident uncertainty estimates and is prone to several optimization challenges such as posterior collapse, especially when coupled with powerful decoders (Blei et al., 2017). Additionally, training the model parameters currently requires back-propagation through an SDE solver, which can be computationally expensive in terms of both time and memory requirements. Potential solutions could lie in emerging simulation-free approaches for training (neural) stochastic differential equations (Course & Nair, 2024; Zhang et al., 2024). Future work could also explore comparing different dynamical systems beyond coupled oscillators and focus on extending the current approach to model neural data from several subjects as well as other recording modalities.
Acknowledgements
This work is supported via the Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research programme Gravitation which is (partly) financed by the Dutch Research Council (NWO).
Appendix
Appendix A Related Work
Latent variable models in neuroscience
Our work builds on previous latent variable models which have been proposed to infer low-dimensional latent dynamics from single-trial data (Sahani, 1999; Yu et al., 2008; Macke et al., 2011; Wu et al., 2017; Pandarinath et al., 2018; Hurwitz et al., 2021b; Schimel et al., 2021; Kim et al., 2021). Most related to our work, is recent work by Hu et al. (2024) which leverage SDEs to represent latent neural dynamics. In their work, the authors developed a novel Gaussian process kernel function that defines a smooth, locally linear prior on dynamics, enabling it to act as an easily interpetable recurrent linear switching dynamical systems while providing posterior uncertainty estimates. A key advantage of our employed approach is the scalability and flexibility which is limited by the use of Gaussian processes. Relevant recent work also by Pals et al. (2024) extend low-rank RNNs to a stochastic setting with additive process noise. The authors use a specific form of the RNN which enables finding fixed-points effectively after training.
Scalable inference of latent SDEs
Inferring the state and learning the parameters of a latent SDE from (high-dimensional) noisy observations is a challenging problem which appears through out science and engineering. With the advent of neural differential equations (Chen et al., 2018; Kidger, 2022), there has been a growing interest in developing scalable methods for fitting latent neural SDEs to data. Seminal work by Tzen & Raginsky (2019a, b) established the theoretical foundation for training generative Neural SDEs via variational inference through the lens of stochastic control. Subsequent research has focused on addressing key computational challenges, including improved memory efficiency and extension to the time domain (Li et al., 2020), enhanced numerical stability (Zeng et al., 2023; Oh et al., 2024), and computational efficiency (Course & Nair, 2024). Our work builds upon the formulation by Li et al. (2020), extending it to controlled and multi-modal settings in neuroscience. This specific formulation enables preserving the expressivity of the generative SDE through state- and input-dependent diffusion on arbitrary spaces.
Modeling neural dynamics with coupled oscillators
Coupled oscillator models have a rich history in theoretical neuroscience, dating back to seminal work by Kuramoto (1975) and Wilson & Cowan (1972). These models have proven particularly effective in capturing rhythmic neural activity and investigating synchronization phenomena in neural circuits (Breakspear et al., 2010; Ermentrout & Terman, 2010). Of particular relevance to our work are modern approaches that combine oscillator models with machine learning techniques (Abrevaya et al., 2021, 2023; Bandyopadhyay et al., 2023) to learn dynamical systems that can reproduce complex neural trajectories while maintaining interpretable structure. Our approach builds upon these foundations by introducing a flexible SDE framework that can naturally capture oscillatory dynamics while accounting for process noise and external inputs. This bridges the gap between mechanistic oscillator models and data-driven approaches, allowing for both accurate trajectory prediction and meaningful scientific interpretation.
Appendix B Learning the parameters of the generative model and inferring latent paths
Recall that our goal is to learn a generative model via the following continuous-time state-space model:
We leverage variational inference to train the model parameters in a tractable and scalable manner. The SDE induces a path measure , which together with the initial distribution implicitly define our prior distribution over continuous paths. To define the posterior distribution, we consider another continuous-time stochastic process defined via the following SDE:
This SDE induces a measure on the path space which together with the initial distribution implicitly define our posterior distribution over continuous paths. The choice of using a similar diffusion function allows us to derive a tractable relationship between the two path measures induced via the two SDEs using Girsanov’s theorem. According to Girsanov’s theorem, the Radon-Nikodym derivative between these measures is given by:
where . We can thus derive a finite Kullback–Leibler divergence between the two path measures as:
The evidence lower bound (ELBO) can then be written as:
This optimization can be performed using stochastic gradient descent, where the path integrals are approximated using numerical integration schemes such as the Euler-Maruyama method, and the expectations are estimated using Monte Carlo sampling.
Appendix C Implementation details
In the following, we summarize the general implementation details of our latent SDE. The models were trained to minimize ELBO on the training set by standard stochastic gradient backprogation through the computational graph. This includes the numerical solver. That is, the intermediate steps are saved in memory during the forward pass and used to update the parameters during the backward pass. While this scales to in terms of memory requirement where is the number of steps in the numerical solver compared to of the adjoint method, we found that this leads to more stable training in our experiments. We used an Euler-Maruyama solver for all our experiments. We found that setting a step size at 2 times the bin width of the spikes in our empirical experiments did not significantly influence the performance and helped improved the training speed and memory requirements.
Notably, we found that cyclic annealing the KL-divergence term in our ELBO significantly improved training stability. We follow the approach by Fu et al. (2019) and use a linear with 4 cycles with a linearly increasing value from 0 to 10 for half the cycle and then stays constant for the second half, until the cycle resets. In experiments which include reconstructing both neural and behavioral data, we use a constant weighting for each during training depending on the dataset. Our experiments were carried out on a NVIDIA A100-PCIE-40GB. Table 2 lists the main components of our latent SDE and their hyperparameters used across all empirical experiments.
| Component | Hyperparameter | Value |
| General | Batch size | 128 |
| Optimizer | Adam | |
| Learning rate | 1e-3 | |
| Max n_epochs | 2000 | |
| Patience | 20 | |
| MC samples (training) | 1 | |
| MC samples (prediction) | 30 | |
| KL annealing scheduler | Cyclic | |
| Differential equation solver | Solver | Euler-Maruyama |
| Step size | ||
| Backpropagation | Discrete-adjoint | |
| (discretize-then-optimize) |
Appendix D Architecture
We summarize below the objective and parameters of each module in our proposed framework, detailing the configurations used in our empirical experiments.
Observation Encoder
The observation encoder serves two purposes: (1) determining the parameters of the posterior distributions for the initial state, and (2) generating a lower-dimensional time-series representation (context) to condition the augmented SDE on the data. This is accomplished through two parallel components: the initialization network and the context network. The initialization network processes the first 30The context network processes the complete time-series of observations through a 2-layer LSTM, producing a time-series of length 64. This context vector undergoes linear interpolation to create a continuous-time representation for the augmented SDE.
Stimulus Encoder
This module creates a lower-dimensional continuous representation of the stimulus. For low-dimensional stimuli (where ), we employ an identity transformation. We include time as an additional input channel. In scenarios without stimulus information (such as the maze reach experiment), we set . After encoding, we use linear interpolation to obtain a continuous-time representation of the stimulus.
Generative SDE
The generative SDE defines our dynamics by mapping the current state and encoded stimulus to the change in the next state. The drift and diffusion structure is flexible and should incorporate prior knowledge when available. In this paper, we present a model combining coupled oscillators with neural networks to represent the SDE’s drift vector field. We also evaluate a neural SDE variant using a state size of 16 and a 2-layer MLP with hidden layer size 64 and tanh activation functions. The diffusion uses a sparse connection layer with non-zero diagonal parameters only, which is shared with the augmented SDE.
Augmented SDE
The augmented SDE induces a path measure on continuous latent paths, implicitly defining our approximate posterior distribution. This neural SDE processes a concatenated vector of the state, encoded stimulus, and context at time to determine the change in the next state. Our implementation uses a 2-layer MLP with hidden size 64 and tanh activation functions for the drift vector field. The diffusion function is identical to that of the generative SDE.
Decoder
The decoder transforms latent states into observations. For systems with both neural and behavioral observations, we implement parallel neural and behavioral decoders. The neural decoder consists of an MLP with one hidden layer of size 128 and soft-plus activations. For neural populations, we employ similar MLPs, each with output size matching the neuron count in its corresponding population. These outputs’ exponents define the rate parameter of a time-inhomogeneous Poisson distribution. The behavioral decoder uses two parallel linear transformations on the latent states to generate the mean and squared variance of a Gaussian distribution at each time step, defining the predicted behavior distribution.

### v2

@@
-Generative Modeling of Neural Dynamics via Latent Stochastic Differential Equations
+Generative Modeling of Neural Dynamics via Latent Stochastic Differential Equations Ahmed ElGazzar1*, Marcel van Gerven1
 Abstract
-We propose a probabilistic framework for developing computational models of biological neural systems.
-In this framework, physiological recordings are viewed as discrete-time partial observations of an underlying continuous-time stochastic dynamical system which implements computations through its state evolution.
-To model this dynamical system, we employ a system of coupled stochastic differential equations with differentiable drift and diffusion functions and use variational inference to infer its states and parameters.
-This formulation enables seamless integration of existing mathematical models in the literature, neural networks, or a hybrid of both to learn and compare different models.
-We demonstrate this in our framework by developing a generative model that combines coupled oscillators with neural networks to capture latent population dynamics from single-cell recordings.
-Evaluation across three neuroscience datasets spanning different species, brain regions, and behavioral tasks show that these hybrid models achieve competitive performance in predicting stimulus-evoked neural and behavioral responses compared to sophisticated black-box approaches while requiring an order of magnitude fewer parameters, providing uncertainty estimates, and offering a natural language for interpretation.
-1 Introduction
+Understanding how neural circuits implement computation remains a central challenge in neuroscience, requiring models that can unite mechanistic interpretability with the complexity of real neural data.
+Current approaches face a fundamental trade-off: mechanistic biophysical and phenomenological models are difficult to fit to single-trial data, while flexible machine learning models lack interpretability and theoretical grounding.
+We propose a probabilistic framework that addresses this challenge by combining domain-specific mathematical models with neural networks through stochastic differential equations (SDEs).
+In this framework, physiological recordings are viewed as discrete-time partial observations of an underlying continuous-time stochastic dynamical system that implements computations through its state evolution.
+We employ coupled SDEs with differentiable drift and diffusion functions and use variational inference to infer system states and parameters.
+This formulation enables seamless integration of existing mathematical models from the literature, neural networks, or hybrid combinations to learn and compare different modeling approaches.
+We demonstrate this framework by developing a generative model that combines coupled oscillators with neural networks to capture latent population dynamics from single-cell recordings.
+Evaluation across three neuroscience datasets spanning different species, brain regions, and behavioral tasks shows that these hybrid models achieve competitive performance in predicting stimulus-evoked neural and behavioral responses compared to sophisticated black-box approaches, while requiring an order of magnitude fewer parameters, providing principled uncertainty estimates, and offering interpretable mathematical descriptions of neural dynamics.
+Introduction
 Unlike many scientific fields, neuroscience lacks a unified theoretical framework that bridges micro-scale mechanisms with macro-scale observations.
-That is, while the biophysical processes underlying individual neuron spiking have been well-understood for decades (Hodgkin & Huxley, 1952), the computational principles governing collective neural activity that give rise to cognition and behavior remain elusive.
-Dynamical systems theory provides a promising framework to address this gap (Durstewitz et al., 2023).
-Within this formalism, the brain is viewed as a complex dynamical system where neural activity patterns evolve through a high-dimensional state space according to well-defined governing equations, providing a mathematical language for understanding how distributed neural circuits implement computation and generate behavior (Van Gelder, 1998; Izhikevich, 2007; Deco et al., 2008; Breakspear, 2017; Favela, 2021).
-This perspective has led to diverse mathematical models, ranging from mean-field approximations of biophysical models that preserve mechanistic interpretability (Wilson & Cowan, 1972; Jirsa & Haken, 1996; Robinson et al., 1997), to phenomenological models inspired by statistical physics that capture emergent collective behavior (Amari, 1977; Hopfield, 1982; Sompolinsky et al., 1988).
-However, while these models have provided key insights into various neural phenomena (Goldman-Rakic, 1995; Wang, 2002; Fries, 2005), fitting them to single-trial neural data has proven challenging (Urai et al., 2022).
+That is, while the biophysical processes underlying individual neuron spiking have been well-understood for decades [33], the computational principles governing collective neural activity that give rise to cognition and behavior remain elusive.
+Dynamical systems theory provides a promising framework to address this gap [60, 18].
+Within this formalism, the brain is viewed as a complex dynamical system, where neural activity patterns evolve through a high-dimensional state space according to well-defined governing equations, providing a mathematical language for understanding how distributed neural circuits implement computation and generate behavior [78, 38, 15, 9, 23].
+This perspective has led to diverse mathematical models, ranging from mean-field approximations of biophysical models that preserve mechanistic interpretability [83, 39, 63] to phenomenological models inspired by statistical physics that capture emergent collective behavior [3, 34, 70].
+However, while these models have provided key insights into various neural phenomena [31, 82, 25], fitting them to single-trial neural data has proven challenging [76].
 In recent years, latent variable models have emerged as a powerful alternative.
-These models represent high-dimensional neural activity in terms of low-dimensional latent states evolving according to Markovian dynamics (Paninski et al., 2010; Hurwitz et al., 2021a).
-Neural computations are thus assumed to be implemented through dynamic motifs in a low-dimensional latent space (e.g., attractors, oscillations, bifurcations) (Vyas et al., 2020; Khona & Fiete, 2022; Durstewitz et al., 2023).
-The popularity of this approach stems from converging empirical evidence that neural population activity during simple tasks lives on a low-dimensional manifold (Gao & Ganguli, 2015).
-Coupled with ongoing developments in machine learning (Kingma & Welling, 2013; Fabius & Van Amersfoort, 2014; Rubanova et al., 2019), latent variable models are becoming a standard tool in systems and computational neuroscience (Chen et al., 2018; Glaser et al., 2020; Hurwitz et al., 2021b; Zhou & Wei, 2020; Kim et al., 2021).
+These models represent high-dimensional neural activity in terms of low-dimensional latent states evolving according to Markovian dynamics [58, 36].
+Neural computations are thus assumed to be implemented through dynamic motifs in a low-dimensional latent space (e.g., attractors, oscillations, bifurcations) [81, 41, 18].
+The popularity of this approach stems from converging empirical evidence that neural population activity during simple tasks lives on a low-dimensional manifold [28].
+Coupled with ongoing developments in machine learning [45, 22, 65], latent variable models are becoming a standard tool in systems and computational neuroscience [11, 30, 37, 89, 43].
 However, selecting an appropriate model for representing latent neural dynamics presents a significant challenge.
@@
 Furthermore, model predictions should account for various sources of uncertainty typically present in such modeling endeavors, such as measurement noise, process noise, and model uncertainty.
-Finally, to utilize the inferred model for online applications as in brain computer interfaces (BCIs), there should be an efficient tractable method for sampling from the model in real-time.
-We posit that recent developments in scientific machine learning, where domain-specific models — often expressed as differential equations — are combined with neural networks and scalable optimization techniques provide an untapped potential to address these challenges (Raissi et al., 2019; Rackauckas et al., 2020; Lai et al., 2021; Karniadakis et al., 2021).
-Additionally, ongoing innovations in training (neural) stochastic differential equations (SDEs) (Tzen & Raginsky, 2019a; Li et al., 2020; Zeng et al., 2023; Course & Nair, 2024), provide an opportunity to develop scalable probabilistic models in neuroscience, a field wherein the systems in question are inherently stochastic (Laing & Lord, 2009; Rolls & Deco, 2010) and the existing measurement tools only provide a coarse proxy of the system under study (Stevenson & Kording, 2011; Urai et al., 2022).
-Contributions
-We develop a general framework for modeling neural dynamics using latent SDEs with differentiable state- and input-dependent drift and diffusion functions.
+Finally, to utilize the inferred model for online applications, as in brain-computer interfaces, there should be an efficient tractable method for sampling from the model in real time.
+We posit that recent developments in scientific machine learning – where domain-specific models, typically expressed using differential equations, are combined with neural networks and scalable optimization techniques – provide an untapped potential to address these challenges [62, 61, 47, 40].
+Additionally, ongoing innovations in training (neural) stochastic differential equations (SDEs) [74, 49, 87, 14], provide an opportunity to develop scalable probabilistic models in neuroscience, a field wherein the systems in question are inherently stochastic [48, 64] and the existing measurement tools only provide a coarse proxy of the system under study [72, 76].
+To make these advances in scientific machine learning available to the neuroscience community, we developed a general framework for modeling neural dynamics using latent SDEs with differentiable state- and input-dependent drift and diffusion functions.
 Within this framework, we introduce a specific generative model, combining coupled oscillators with neural networks, to capture neural population dynamics.
@@
 Finally, our results suggest that data-driven coupled oscillators could provide an expressive yet interpretable framework for understanding neural dynamics, while maintaining computational tractability.
-2 Framework
-Notations
-Let denote a continuous path of latent neural states, which we observe through neural recordings (e.g., spike counts, fMRI time-series) denoted by , where each is observed at time .
-Additionally, we may observe behavioral responses (e.g., movement trajectories, choice responses) represented as , where each .
-In some settings, these observations are driven by external stimuli or task variables (e.g., visual stimuli, auditory cues) denoted by , where each .
-A single trial is thus defined as , where , and/or could be missing.
+Related work
+Latent variable models in neuroscience
+Our work builds on previous latent variable models which have been proposed to infer low-dimensional latent dynamics from single-trial data [68, 86, 50, 85, 57, 16, 37, 69, 43, 44].
+Most related to our work, is Gaussian process switching linear dynamical system [2] which leverage SDEs to represent latent neural dynamics extending previous work proposed by Duncker et.
+al [16].
+In their work, the authors developed a novel Gaussian process kernel function that defines a smooth, locally linear prior on dynamics, yielding easily interpretable recurrent linear switching dynamical systems while providing posterior uncertainty estimates.
+Compared to Gaussian process-based methods, our approach provides flexibility by allowing the modeler to choose any dynamical system as the prior, enabling systematic model comparison across a wide range of existing neural models.
+Additionally, our approach is not limited to a small number of latent state parameters as is often the case with GP-based approaches, enabling greater flexibility in model capacity.
+Finally, unlike [2], the SDEs employed in our approach are not restricted to additive process noise, but can model multiplicative (state- and input-dependent) process noise, enabling greater flexibility and model expressivity.
+Other relevant recent work extends low-rank RNNs to a stochastic setting with additive process noise [56].
+Here, the authors use a specific form of the RNN which enables finding fixed-points effectively after training.
+Scalable inference of latent SDEs
+Inferring the state and learning the parameters of a latent SDE from (high-dimensional) noisy observations is a challenging problem which appears throughout science and engineering.
+With the advent of neural differential equations [11, 42], there has been a growing interest in developing scalable methods for fitting latent neural SDEs to data.
+Seminal work by [74, 75] established the theoretical foundation for training generative neural SDEs via variational inference through the lens of stochastic control.
+Subsequent research has focused on addressing key computational challenges, including improved memory efficiency and extension to the time domain [49], enhanced numerical stability [87, 55], and computational efficiency [14].
+Our work builds upon the formulation by [49], extending it to controlled and multi-modal settings in neuroscience.
+This specific formulation enables preserving the expressivity of the generative SDE through state- and input-dependent diffusion on arbitrary spaces.
+Modeling neural dynamics with coupled oscillators
+Coupled oscillator models have a rich history in theoretical neuroscience, dating back to seminal work by [46] and [83].
+These models have proven particularly effective in capturing rhythmic neural activity and investigating synchronization phenomena in neural circuits [8, 21].
+Of particular relevance to our work are modern approaches that combine oscillator models with machine learning techniques [1, 2, 4] to learn dynamical systems that can reproduce complex neural trajectories while maintaining interpretable structure.
+Our approach builds upon these foundations by introducing a flexible SDE framework that can naturally capture oscillatory dynamics while accounting for process noise and external inputs.
+This bridges the gap between mechanistic oscillator models and data-driven approaches, allowing for both accurate trajectory prediction and meaningful scientific interpretation.
+Methods
+Notation
+Let denote a continuous path of latent neural states, which we observe through neural recordings (e.g., spike counts, fMRI time-series) denoted by with is observed at time .
+Additionally, we may observe behavioral responses (e.g., movement trajectories, choice responses) represented as with .
+In some settings, these observations are driven by external stimuli or task variables (e.g., visual stimuli, auditory cues) denoted by with .
+A single trial is thus defined as , where , and/or could be (partially) missing.
 Given a dataset consisting of trials, our objectives are to: (1) infer the underlying neural dynamics that generated the observations, represented by the posterior distribution , (2) learn a generative model capable of predicting stimulus-evoked responses, represented by the joint distribution , and (3) extract interpretable dynamical features that provide insights into neural computations.
-2.1 Generative model
+Generative model
 We model latent neural dynamics as a continuous-time stochastic process, specified as the solution to the following Itô SDE:
@@
 This SDE induces a path measure on the space of continuous paths which implicitly define the distribution .
-Samples from this distribution are obtained by solving the SDE with different realizations of and :
+Samples from this distribution for are obtained by solving the SDE with different realizations of and :
 The continuous-time latent states are mapped to discrete-time observations through:
-where and denote the observation models (e.g., Poisson model for spike counts, Gaussian model for continuous measurements), with and being differentiable functions mapping the latent state to the parameters of these distributions.
-2.2 Inference and learning
-The true parameters of our generative model are unknown and need to be learned from the data .
+where and denote the observation models (e.g., Poisson model for discrete measurements, Gaussian model for continuous measurements) for neural and behavioral data, respectively, with and being differentiable functions mapping the latent state to the parameters of these distributions.
+Model estimation
+The true parameters of our generative model are unknown and need to be learned from data .
 Moreover, the exact posterior distribution is intractable to compute.
 Variational inference provides an efficient and scalable approach to tackle this problem.
-However unlike standard latent variable models, the latent state in our model are realizations of a continuous-time stochastic process.
+Unlike standard latent variable models, the latent state in our model are realizations of a continuous-time stochastic process.
 Thus we need to define a parametric distribution over paths rather than random-valued vectors.
-Following (Tzen & Raginsky, 2019a, b; Li et al., 2020), we define the approximate posterior distribution via an augmented neural SDE.
-Specifically, here we define the approximate posterior path as realizations of the following SDE (we omit time indices for ease of notation):
-where and is a continuous path encoding the observations, with being a modality-specific observation encoder.
-We define the distribution of the initial condition as , where and are neural networks, and .
-We further set to prevent over-parameterizing the initial conditions and rendering the dynamics redundant (Zabihi et al., 2019).
-Here is a neural network representing the drift of the augmented SDE.
-Note that while the drift term is different from that of the generative model, the diffusion term is shared.
-As illustrated in (Tzen & Raginsky, 2019a), this specification enables leveraging Girsanov’s theorem (Girsanov, 1960) to obtain a tractable KL divergence between the the generative (prior) and augmented (approximate posterior) distributions:
+Figure 1 provides an overview of our framework.
+Following [74, 75, 49], we define the approximate posterior distribution via an augmented (neural) SDE.
+Specifically, we define the approximate posterior path as realizations of the following SDE (we omit time indices for ease of notation):
+where and is a continuous path encoding the observations which we refer to a context, with being a modality-specific observation encoder.
+We define the distribution of the initial condition as , where and are neural networks, and are discrete samples of the continuous-time context .
+Here, is a neural network representing the drift term of the augmented SDE.
+Note that while the drift term is different from that of the generative model, the diffusion term is shared with the generative SDE.
+As described in [74], this specification enables leveraging Girsanov’s theorem [29] to obtain a tractable KL divergence between the generative (prior) and augmented (approximate posterior) distributions:
 where with and denoting the measures induced by the generative and augmented SDE, respectively.
-Using this we can define an evidence lower bound (ELBO) on the conditional marginal likelihood of the observations:
-The ELBO consists of three main terms: (1) the expected log-likelihood of the observations under the approximate posterior, which includes both neural recordings and behavioral responses, (2) the KL divergence between the initial state distributions, and (3) the path-wise KL divergence between the approximate posterior and prior processes derived via Girsanov’s theorem (see Appendix B for detailed derivation of the loss).
+Using this property, we can define the following evidence lower bound (ELBO) on the conditional marginal likelihood of the observations:
+The ELBO consists of three main terms: (1) the expected log-likelihood of the observations under the approximate posterior, which includes both neural recordings and behavioral responses, (2) the KL divergence between the initial state distributions, and (3) the path-wise KL divergence between the approximate posterior and prior processes derived via Girsanov’s theorem (see Supplementary Information for a detailed derivation of the loss).
 The model parameters and variational parameters can be jointly optimized by maximizing this ELBO using stochastic gradient descent.
-Both the expectation over and the path-wise integral in the KL divergence term are computed using numerical integration schemes suitable for SDEs (see Appendix C for implementation details).
+Both the expectation over and the path-wise integral in the KL divergence term are computed using numerical integration schemes suitable for SDEs (see Supplementary Information for implementation details).
 This formulation enables us to simultaneously learn the parameters of the generative model while performing approximate posterior inference over the latent neural trajectories.
-The learned generative model can then be used for various online and offline downstream tasks including probabilistic prediction of stimulus-evoked responses and extraction of interpretable dynamical features.
-3 Hybrid modeling
-So far we have described a general framework for modeling the latent dynamics of a neural system from empirical observations.
+The learned generative model can then be used for various online and offline downstream tasks, including probabilistic prediction of stimulus-evoked responses and extraction of interpretable dynamical features.
+Biophysics-inspired latent SDE models
+So far, we have described a general framework for modeling the latent dynamics of a neural system from empirical observations.
 This framework allows us to leverage almost arbitrary differentiable functions to model the drift and diffusion terms of an SDE.
 A natural choice would be to use a neural SDE to act as the generative model because of their universal approximation properties.
-However, this comes at the cost of efficiency and interpetability.
+However, this comes at the cost of efficiency and interpretability.
 An alternative option, is to leverage prior models proposed in the literature and use neural networks either to approximate their parameters or to augment their dynamics.
-Such hybrid models can improve sample-efficiency, generalization performance and, are easier to interpret in comparison to models that rely only on neural networks (Rackauckas et al., 2020; ElGazzar & van Gerven, 2024).
+Such hybrid models can improve sample-efficiency and generalization performance, and are easier to interpret in comparison to models that rely only on neural networks [61, 20].
 In this section, we detail a specific instantiation of the framework which combines a well-understood dynamical system and neural networks to obtain an expressive, interpretable and parameter-efficient generative model of neural population dynamics.
-3.1 Stochastic nonlinear coupled oscillators
+Furthermore, we describe how such models can be linked to neurobehavioral data via the specification of appropriate forward models.
+Stochastic nonlinear coupled oscillators
 A well-known class of models that are used to study several biological and physical systems are coupled oscillators.
-Coupled oscillators provide a rich yet simplified language for studying complex interactions in systems exhibiting emergent behavior such as synchronization, pattern formation and phase transitions (Winfree, 1980).
-We focus on a specific instantiation of coupled oscillators (Matthews et al., 1991) described by the following complex-valued ODE:
+Coupled oscillators provide a rich yet simplified language for studying complex interactions in systems exhibiting emergent behavior such as synchronization, pattern formation and phase transitions [84].
+We focus on a specific instantiation of coupled oscillators [53] described by the following complex-valued ordinary differential equation (ODE):
 where represents the position of oscillators in the complex plane, the bifurcation parameters, denotes the natural frequency, and represents all-to-all coupling strength.
@@
 Specifically, we use the position of the oscillator as the latent state in the generative SDE.
-Additionally, we consider the coupling strength as a time-varying input dependent function of the external input .
+Additionally, we consider the coupling strength as a time-varying input-dependent function of the external input .
 Incorporating these updates and writing the model in terms of the real and imaginary parts of the oscillator , we obtain the following system of SDEs:
-where , , are neural networks representing the input dependent connectivity and the diffusion terms for the real and imaginary states respectively.
+where represents the state of coupled oscillators and , , are neural networks representing the input dependent connectivity and the diffusion terms for the real and imaginary states respectively.
 The trainable parameters of the generative SDE are thus the parameters of the neural networks as well as and .
-We refer to this dynamics model as coupled oscillator SDE (CO-SDE).
-3.2 Modeling multiple interacting populations
-Often, we have access to neural activity recorded from neurons across multiple brain regions, simultaneously.
-Given access to recordings from brain regions, we can extend the generative model by factorizing the latent state where the number of oscillators can vary depending on the population size and complexity.
-The observed activity of a population is then
+We refer to this dynamics model as a coupled-oscillator SDE (CO-SDE).
+Modeling neurobehavioral observations
+Often, we have access to observational data, which can come in the form of neural activity recorded from multiple brain regions or in the form of behavioral measurements.
+To inform the latent SDE model, we need to condition the generative model on such observational data.
+Given access to neural activity recorded from brain regions, we can extend the generative model by factorizing the latent state , where the number of oscillators can vary depending on the population size and complexity.
+Assuming that the neural data comes in the form of population firing rates, we model the observed activity of a population as
 where is a vector of rate parameters which depends on a population-specific transformation that maps the latent state to the firing rates of the population.
-4 Experiments
-4.1 Simulated data
-We first sought to evaluate the efficacy of our framework on modeling a partially observed neural system with different levels of process noise.
+For behavioral observations, in case of continuous-valued data (e.g. arm position, pupil diameter), we assume a Gaussian measurement model:
+where and are neural networks that map the latent state to the mean and variance parameters of the observation distribution, respectively.
+Finally, for discrete-valued behavioral data (e.g., choice response), we employ a categorical observation model:
+where is a neural network followed by a softmax activation that maps the latent state to a probability vector over discrete categories.
+Note that neural and behavioral data can be easily combined by using the described forward models in parallel.
+This provides a natural solution to the data fusion problem and provides a principled way to integrate multi-modal data in our analyses.
+Results
+Validation on simulated data
+We first sought to evaluate the efficacy of our framework for modeling a partially observed neural system with different levels of process noise.
 To this end, we generated spiking data from the following system:
 where describes the population average firing rate of three neural populations driven by a sinusoidal input with randomly sampled frequency and phase.
-and describe the recurrent connectivity and input interaction matrices, respectively.
+The system matrix and input matrix describe the recurrent connectivity and input dependence, respectively.
 The time scale of the network dynamics is given by and describes the magnitude of process noise in the system.
 The system output is given by representing the observed spike counts of 50 neurons generated by sampling from an inhomogeneous Poisson process with rates computed by affine transformation of the population average activity via the output matrix .
-We fit our latent SDE model on the spiking data by maximizing the Poisson-log likelihood of the observations given predicted rate and evaluated performance using bits per spike (bps), a metric that quantifies how well the predicted firing rates explain the observed spike patterns (Pei et al., 2021).
+We fit our latent SDE model on the spiking data by maximizing the Poisson-log likelihood of the observations given predicted rate and evaluated performance using bits per spike (bps), a metric that quantifies how well the predicted firing rates explain the observed spike patterns [59].
 For our experiments, we varied the noise level to test the robustness of the model.
-We used a latent neural SDE with varying sizes of latent dimension and compared against two baselines: a latent ODE (Rubanova et al., 2019) and LFADS (Pandarinath et al., 2018).
-As demonstrated in Figure 2a, the latent SDE model can better explain the data variance when using fewer latent dimensions compared to a latent ODE model.
+We used a latent neural SDE with varying sizes of latent dimension and compared against two baselines: a latent ODE [65] and LFADS [57] as these have gained popularity in the field.
+As demonstrated in Fig.
+2a, the latent SDE model can better explain the data variance when using fewer latent dimensions compared to a latent ODE model.
 Increasing the latent dimensions beyond the true state dimensions does not significantly improve the model fit unlike the ODE-based model which stagnates at a lower absolute value and higher latent dimension.
-Additionally, as shown in Figure 2b, the latent SDE model is more robust to process noise compared to latent ODE and LFADS models.
-These results suggest that latent SDEs can be more reliable to model real-world data when the true state dimension is unknown and process noise can emerge from unobserved interactions.
+Additionally, as shown in Fig.
+2b, the latent SDE model is more robust to process noise compared to latent ODE and LFADS models.
+These results suggest that latent SDEs can be more reliable when modeling real-world data when the true state dimension is unknown and process noise can emerge from unobserved interactions.
 In this setting, a probabilistic initial state with deterministic dynamics, as in the case of latent ODEs and LFADS, may not be sufficient.
+Figures 2c and 2d show the model predictions for the spike counts and the underlying firing rates, respectively.
+Figure 2e demonstrates the phase portrait of the inferred dynamical system compared to the ground truth simulated dynamical system.
+Validation on empirical data
 | Datasets | ||||||
@@
 | Neural ODE | 4880 | 0.221 ± 0.005 | 0.86 ± 0.012 | 0.15 ± 0.01 | 0.59 ± 0.025 | 0.316 ± 0.008 |
-| Coupled Oscillator ODE | 465 | 0.221 ± 0.006 | 0.85 ± 0.015 | 0.21 ± 0.009 | 0.61 ± 0.02 | 0.309 ± 0.007 |
+| CO-ODE | 465 | 0.221 ± 0.006 | 0.85 ± 0.015 | 0.21 ± 0.009 | 0.61 ± 0.02 | 0.309 ± 0.007 |
+| GP-SLDS | NA | 0.201 ± 0.004 | NA | 0.08 ± 0.01 | NA | 0.322 ± 0.004 |
+| Linear SDE | 304 | 0.147 ± 0.01 | 0.81 ± 0.02 | 0.18 ± 0.012 | 0.63 ± 0.02 | 0.207 ± 0.01 |
 | Neural SDE | 4944 | 0.243 ± 0.007 | 0.86 ± 0.015 | 0.18 ± 0.012 | 0.70 ± 0.02 | 0.317 ± 0.009 |
-| Coupled Oscillator SDE | 526 | 0.238 ± 0.008 | 0.86 ± 0.018 | 0.22 ± 0.011 | 0.67 ± 0.025 | 0.314 ± 0.01 |
-4.2 Empirical data
-4.2.1 Datasets & baselines
+| CO-SDE | 526 | 0.238 ± 0.008 | 0.86 ± 0.018 | 0.22 ± 0.011 | 0.67 ± 0.025 | 0.314 ± 0.01 |
+Datasets & baselines
 We evaluated our framework across three neuroscience datasets spanning different brain regions and behavioral tasks.
-Perturbed reach task In this experiment, a monkey is presented with a target location and tasked with moving the cursor from the center to the target location by controlling a manipulandum (Chowdhury et al., 2020) (Figure 3a).
+Perturbed reach task
+In this experiment, a monkey is presented with a target location and tasked with moving the cursor from the center to the target location by controlling a manipulandum [12] (Fig.
+3a).
 On one portion of the trials, the monkey’s arm was perturbed before the reach by applying resisting forces on the manipulandum.
@@
 The input to the model are the target location, as well as the perturbing force and torques on the manipulandum.
-Visual decision making In this task, a mouse is presented with images of different contrasts and the mouse is tasked with turning a wheel to bring the image with higher contrast to the center of the screen when a go cue is randomly presented (Steinmetz et al., 2019) (Figure 3d).
+At test-time, we provide the first 50 ms of observations before stimulus presentation to the observation encoder to identify the initial condition.
+Visual decision making
+In this task, a mouse is presented with images of different contrasts and the mouse is tasked with turning a wheel to bring the image with higher contrast to the center of the screen when a go cue is randomly presented [71] (Fig.
+3d).
 Neural activity is recorded in multiple regions across the brain using Neuropixels probes.
@@
 The input to the model are the contrast levels, the timing of their presentation, and the timing of the go cue.
-Delayed reach task The task is a delayed center-out reach task with barriers, resulting in a variety of straight and curved trajectories (Mante et al., 2013) (Figure 3g).
-Neural activity was recorded from the dorsal premotor (PmD) and primary motor (M1) cortices, and cursor, monkey gaze position, and monkey hand position and velocity are also provided.
+At test-time, we provide the first 50 ms of observations before stimulus presentation to the observation encoder to identify the initial conditions.
+Delayed reach task
+The task is a delayed center-out reach task with barriers, resulting in a variety of straight and curved trajectories [52] (Fig.
+3g).
+Neural activity was recorded from the dorsal pre-motor (PmD) and primary motor (M1) cortices, and cursor, monkey gaze position, and monkey hand position and velocity are also provided.
 Here we train a model to predict neural responses during movement in both areas given the preparatory activity before the go cue.
 No stimulus information is provided to the model during training or prediction.
+At test-time, we provide the first 100 ms of neural observations before the go cue to the observation encoder to identify the initial conditions.
 We compare different latent variable models across the three datasets.
-To enable a fair compariosn, we fix the encoder and decoder architectures across all models while varying the latent sequential model.
-That is, we compare the performance of RNN, GRU, LSTM, neural ODE, and coupled oscillator ODE models with that of our proposed coupled oscillator SDE and a neural SDE model.
+To enable a fair comparison, we fix the encoder and decoder architectures across all models while varying the latent sequential model.
+That is, we compare the performance of RNN, GRU, LSTM, Neural ODE, and a coupled oscillator ODE (CO-ODE) models with that of our proposed CO-SDE and a Neural SDE model.
+We also include a linear SDE model which employs linear drift and diffusion functions.
 All the baseline models had a probabilistic initial state and were trained via variational inference.
-Note that this formulation enables generalizing the baselines to several previous models proposed in literature (Pandarinath et al., 2018; Hurwitz et al., 2021b; Kim et al., 2021; Versteeg et al., 2023) (See Appendix D for details on the model architecture).
-4.2.2 Empirical results
-Table 1 presents 5-fold cross-validation results across the three datasets, evaluating both neural activity prediction (in bits per spike) and behavioral performance (via the coefficient of determination, ).
+Note that this formulation enables generalizing the baselines to several previous models proposed in literature [57, 37, 43, 80].
+We also compare against a recently proposed Gaussian process switching linear dynamical system (GP-SLDS) [2] (See Supplementary Information for details on the baselines).
+Empirical results
+Table 1 presents 5-fold cross-validation results across the three datasets, evaluating both neural activity prediction (in bits per spike) and behavioral performance (via the coefficient of determination, ) (see Supplementary Information for further details on the evaluation metrics).
 First, we observe that our stochastic models (Neural SDE and CO-SDE) on average consistently outperform their deterministic counterparts.
 This suggests that incorporating stochasticity into the latent dynamics is crucial for capturing the variability inherent in neural and behavioral data and that a stochastic initial state might not be sufficient to capture uncertainty in the dynamics.
-Second, we observe that model complexity, as measured by parameter count, does not correlate with performance.
+Second, we find that non-linear latent dynamics are crucial for accurate joint prediction of neural and behavioral responses.
+This is suggested by the performance of the linear SDE baseline which performs competitively to non-linear baselines only on predicting behavioral responses on the perturbed reach task.
+However, its performance is significantly lower for predicting neural responses across all three tasks.
+This suggests that while linear dynamics may suffice for capturing certain behavioral regularities, nonlinear models are necessary for accurate neural population modeling at the level of single-neuron.
+Third, we observe that model complexity, as measured by parameter count, does not correlate with performance.
 The coupled oscillator models achieve competitive or superior performance while using only 465-526 parameters, compared to 33,280 parameters for the LSTM model.
@@
 This also suggests that standard deep learning models may be too over-parameterized to learn the underlying dynamical system during simple behavioral tasks.
-Beyond quantitative results, oscillator-based models provide a direct way to interpret latent neural dynamics.
-We visualize the value of the inferred frequencies of the model after training as well as the value of the coupling strength in response to sample stimuli in Figure 3.
-In the perturbed reach task, where we have data from one somatosensory area, we observe that the connectivity strength increases over time (Figure 3c).
-This is also observed in the delayed reach task where we have recordings from two regions in the motor in the motor cortex (Figure 3k).
-In contrast, on the visual decision making task where we have recordings from multiple regions across the brain the coupling strength decreases over time (Figure 3f).
-This decrease is marked via two significant drops at two timepoints which mark the presentation of the stimulus and the go cue, respectively.
-Additionally, we observe that on average, the inferred natural frequency of the oscillators differ between brain regions.
-Notably, the average frequency of oscillators corresponding to M1 is higher than the average frequency of oscillators corresponding to PMd during the delayed reach task (Figure 3k), which may be linked to distinct functional roles related to action preparation and execution (Thura et al., 2022).
-5 Discussion
+Beyond quantitative performance, oscillator-based models offer interpretable parameterizations that connect to established neuroscience concepts.
+We visualize the inferred natural frequencies and coupling strengths learned by the model in Fig.
+3.
+In both motor tasks —the perturbed reach task with recordings from somatosensory cortex and the delayed reach task with recordings from two motor cortical areas- we observe that coupling strength increases over time during movement preparation and execution (Fig.3c,k).
+This pattern is consistent with theories of neural communication through enhanced synchronization [26], where increasing oscillatory coupling may facilitate information integration across neural populations during coordinated motor output.
+In contrast, the visual decision-making task shows a markedly different pattern (Fig.3f).
+Here, coupling strength decreases over time, with two distinct drops coinciding with the stimulus presentation and go cue.
+This temporal modulation may reflect dynamic routing of information[26], where coupling is selectively reduced to gate irrelevant signals or reallocate processing resources as task demands shift from sensory encoding to decision formation and action selection.
+We also observe systematic differences in natural frequencies across brain regions.
+Notably, oscillators corresponding to M1 exhibit higher average frequencies than those in PMd during the delayed reach task (Fig.3h).
+This frequency gradient is consistent with the hierarchical organization of motor cortex, where M1 operates at faster timescales than PMd, which is more involved in movement planning [73, 13].
+Such frequency differences may reflect distinct computational roles: higher frequencies in M1 may support rapid sensorimotor transformations required for real-time motor control, while lower frequencies in PMd may be suited for integrating information over longer timescales during motor preparation.
+Important note on interpretation: While these learned parameters show relative consistency across model initializations and align with neuroscience literature, we emphasize that our method does not provide theoretical guarantees for parameter identifiability—a well-known challenge in nonlinear dynamical systems [54].
+Therefore, we present these findings as hypothesis-generating observations that suggest our model captures functionally meaningful structure, rather than as definitive claims about the underlying neural mechanisms.
+The value of these parameterizations lies in providing neuroscience-motivated, interpretable descriptions of population dynamics that can guide future experimental investigations.
+Discussion
 This work introduces a probabilistic framework for modeling neural dynamics that combines the expressiveness of neural networks with the interpretability of mechanistic or phenomenological models.
@@
 First, explicitly modeling probabilistic dynamics through SDEs consistently improves model performance compared to deterministic alternatives, even when the latter incorporate uncertainty through probabilistic initial conditions.
-This suggests that neural variability stems not just from uncertain initial states but from inherently stochastic dynamics (Laing & Lord, 2009).
+This suggests that neural variability stems not just from uncertain initial states but from inherently stochastic dynamics [48].
 The superior performance of SDE-based models in high-noise regimes as shown in the simulations results further supports this conclusion.
-Additional simulation results also show that using stochastic dynamics can improve model fit while relying on lower-dimensional latent states; an observation that is also found in stochastic low-rank RNNs (Pals et al., 2024).
+Additional simulation results also show that using stochastic dynamics can improve model fit while relying on lower-dimensional latent states; an observation that is also found in stochastic low-rank RNNs [56].
 Second, our hybrid coupled oscillator model achieves competitive or superior performance while requiring an order of magnitude fewer parameters than standard deep learning approaches.
-This parameter efficiency suggests that incorporating appropriate inductive biases through mechanistic components can effectively constrain the model space (Rackauckas et al., 2020; ElGazzar & van Gerven, 2024).
-The success of coupled oscillators in particular hints at their potential as a natural description language for neural population dynamics, consistent with existing theoretical work on neural oscillations (Buzsaki & Draguhn, 2004; Breakspear et al., 2010; Churchland et al., 2012; Bick et al., 2020).
-This also supports the potential of coupled oscillators as an expressive and efficient tool to model arbitrary time-series data, as seen in emerging machine learning research (Rusch & Mishra, 2020; Effenberger et al., 2022; Rusch & Rus, 2024).
-With that said, there are a number of limitations to the proposed approach which merits consideration.
-First, the framework relies on variational inference for training the model parameters.
-This could be suboptimal as variational inference is known to provide overconfident uncertainty estimates and is prone to several optimization challenges such as posterior collapse, especially when coupled with powerful decoders (Blei et al., 2017).
-Additionally, training the model parameters currently requires back-propagation through an SDE solver, which can be computationally expensive in terms of both time and memory requirements.
-Potential solutions could lie in emerging simulation-free approaches for training (neural) stochastic differential equations (Course & Nair, 2024; Zhang et al., 2024).
-Future work could also explore comparing different dynamical systems beyond coupled oscillators and focus on extending the current approach to model neural data from several subjects as well as other recording modalities.
+This parameter efficiency suggests that incorporating appropriate inductive biases through mechanistic components can effectively constrain the model space [61, 20].
+The success of coupled oscillators in particular hints at their potential as a natural description language for neural population dynamics, consistent with existing theoretical work on neural oscillations [10, 8, 13, 6].
+This also supports the potential of coupled oscillators as an expressive and efficient tool to model arbitrary time-series data, as seen in emerging machine learning research [66, 19, 67].
+Future work could explore in depth the connection between features of inferred coupled oscillators (e.g.
+synchronization, phase-amplitude coupling, etc.) and observed neural and behavioral responses.
+While we focus on coupled oscillator models motivated by motor cortex physiology, our framework readily accommodates diverse dynamical systems including Van der Pol oscillators [77] for modeling limit cycles with different relaxation properties, FitzHugh-Nagumo [24] models for excitable systems, attractor-based models for decision-making circuits [82], or predator-prey dynamics for competitive interactions [27].
+The choice of dynamical model should be guided by domain knowledge and scientific hypotheses about the neural system.
+This flexibility enables researchers to test specific mechanistic hypotheses while maintaining the benefits of probabilistic inference and uncertainty quantification.
+Future work could systematically compare these alternative models on the same datasets to identify which dynamical structures best capture different neural computations.
+Alternatively, the symbolic expressions that define the latent dynamics could be inferred using black-box optimization techniques such as differential genetic programming [79].
+With that said, there are a number of limitations to the proposed framework which merits consideration.
+First, similar to most non-linear latent variable models in the literature, our approach does not guarantee parameter identifiability.
+Rather, our approach enables interpreting the system dynamics as a function of the behavior of the proposed system dynamics.
+Further work is required to ensure identifiability of the parameters of the latent dynamical system to enable robust comparison of different models [17, 32].
+Here, advances in epistemic AI that focus on quantifying parameter uncertainty are of particular relevance [51].
+Another limitation is that our framework relies on variational inference for training the model parameters.
+This could be suboptimal as variational inference is known to provide overconfident uncertainty estimates and is prone to several optimization challenges such as posterior collapse, especially when coupled with powerful decoders [7].
+That is, the model can rely on expressive decoders to model the data, rendering latent variables redundant.
+A potential solution is to utilize invertible decoders [89, 80] to guarantee that the model relies on latent dynamics to model neural and behavioral observations.
+Furthermore, training the model parameters currently requires back-propagation through an SDE solver, which can be computationally expensive in terms of both time and memory requirements.
+Potential solutions could lie in emerging simulation-free approaches for training (neural) stochastic differential equations [14, 88, 5].
+In this work, we provided a first demonstration of the usefulness of developing biophysics-inspired latent SDE models.
+Future work could explore comparing different dynamical systems beyond coupled oscillators and focus on extending the current approach to model neural data from several subjects as well as other recording modalities.
+The use of these models for neural system identification may provide new insights into neural information processing.
+Furthermore, their interpretability and expressiveness may pave the way for new applications in clinical neuroscience.
 Acknowledgements
-This work is supported via the Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research programme Gravitation which is (partly) financed by the Dutch Research Council (NWO).
-Appendix
-Appendix A Related Work
-Latent variable models in neuroscience
-Our work builds on previous latent variable models which have been proposed to infer low-dimensional latent dynamics from single-trial data (Sahani, 1999; Yu et al., 2008; Macke et al., 2011; Wu et al., 2017; Pandarinath et al., 2018; Hurwitz et al., 2021b; Schimel et al., 2021; Kim et al., 2021).
-Most related to our work, is recent work by Hu et al.
-(2024) which leverage SDEs to represent latent neural dynamics.
-In their work, the authors developed a novel Gaussian process kernel function that defines a smooth, locally linear prior on dynamics, enabling it to act as an easily interpetable recurrent linear switching dynamical systems while providing posterior uncertainty estimates.
-A key advantage of our employed approach is the scalability and flexibility which is limited by the use of Gaussian processes.
-Relevant recent work also by Pals et al.
-(2024) extend low-rank RNNs to a stochastic setting with additive process noise.
-The authors use a specific form of the RNN which enables finding fixed-points effectively after training.
-Scalable inference of latent SDEs
-Inferring the state and learning the parameters of a latent SDE from (high-dimensional) noisy observations is a challenging problem which appears through out science and engineering.
-With the advent of neural differential equations (Chen et al., 2018; Kidger, 2022), there has been a growing interest in developing scalable methods for fitting latent neural SDEs to data.
-Seminal work by Tzen & Raginsky (2019a, b) established the theoretical foundation for training generative Neural SDEs via variational inference through the lens of stochastic control.
-Subsequent research has focused on addressing key computational challenges, including improved memory efficiency and extension to the time domain (Li et al., 2020), enhanced numerical stability (Zeng et al., 2023; Oh et al., 2024), and computational efficiency (Course & Nair, 2024).
-Our work builds upon the formulation by Li et al.
-(2020), extending it to controlled and multi-modal settings in neuroscience.
-This specific formulation enables preserving the expressivity of the generative SDE through state- and input-dependent diffusion on arbitrary spaces.
-Modeling neural dynamics with coupled oscillators
-Coupled oscillator models have a rich history in theoretical neuroscience, dating back to seminal work by Kuramoto (1975) and Wilson & Cowan (1972).
-These models have proven particularly effective in capturing rhythmic neural activity and investigating synchronization phenomena in neural circuits (Breakspear et al., 2010; Ermentrout & Terman, 2010).
-Of particular relevance to our work are modern approaches that combine oscillator models with machine learning techniques (Abrevaya et al., 2021, 2023; Bandyopadhyay et al., 2023) to learn dynamical systems that can reproduce complex neural trajectories while maintaining interpretable structure.
-Our approach builds upon these foundations by introducing a flexible SDE framework that can naturally capture oscillatory dynamics while accounting for process noise and external inputs.
-This bridges the gap between mechanistic oscillator models and data-driven approaches, allowing for both accurate trajectory prediction and meaningful scientific interpretation.
-Appendix B Learning the parameters of the generative model and inferring latent paths
+This publication is part of the project Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research programme Gravitation which is (partly) financed by the Dutch Research Council (NWO).
+Supplementary Information
+1 Learning the parameters of the generative model and inferring latent paths
 Recall that our goal is to learn a generative model via the following continuous-time state-space model:
@@
 This optimization can be performed using stochastic gradient descent, where the path integrals are approximated using numerical integration schemes such as the Euler-Maruyama method, and the expectations are estimated using Monte Carlo sampling.
-Appendix C Implementation details
+2 Implementation details
 In the following, we summarize the general implementation details of our latent SDE.
@@
 We found that setting a step size at 2 times the bin width of the spikes in our empirical experiments did not significantly influence the performance and helped improved the training speed and memory requirements.
-Notably, we found that cyclic annealing the KL-divergence term in our ELBO significantly improved training stability.
-We follow the approach by Fu et al.
-(2019) and use a linear with 4 cycles with a linearly increasing value from 0 to 10 for half the cycle and then stays constant for the second half, until the cycle resets.
+We ran a hyperparameter search for the latent states, which in our oscillatory models (CO-ODE, CO-SDE) represents the number of oscillators .
+We found that the optimal number of oscillators varies across tasks (typically between 8 and 16) for the evalutaed tasks.
+We also found that cyclic annealing the KL-divergence term in our ELBO significantly improved training stability.
+We use a linear with 4 cycles with a linearly increasing value from 0 to 10 for half the cycle and then stays constant for the second half, until the cycle resets.
 In experiments which include reconstructing both neural and behavioral data, we use a constant weighting for each during training depending on the dataset.
@@
 | (discretize-then-optimize) |
-Appendix D Architecture
+3 Architecture
 We summarize below the objective and parameters of each module in our proposed framework, detailing the configurations used in our empirical experiments.
@@
 This is accomplished through two parallel components: the initialization network and the context network.
-The initialization network processes the first 30The context network processes the complete time-series of observations through a 2-layer LSTM, producing a time-series of length 64.
+The initialization network processes the first 30% of the observations using a 2-layer LSTM with hidden size 64, applying it in reverse order.
+The network then transforms the output through two parallel affine transformations into the latent dimension size, with a soft-plus activation applied to the variance component.
+This generates the mean and squared variance of the posterior initial condition.
+The context network processes the complete time-series of observations through a 2-layer LSTM, producing a time-series of length 64.
 This context vector undergoes linear interpolation to create a continuous-time representation for the augmented SDE.
@@
 In this paper, we present a model combining coupled oscillators with neural networks to represent the SDE’s drift vector field.
-We also evaluate a neural SDE variant using a state size of 16 and a 2-layer MLP with hidden layer size 64 and tanh activation functions.
+We also evaluate a neural SDE variant with a 2-layer MLP with hidden layer size 64 and tanh activation functions.
 The diffusion uses a sparse connection layer with non-zero diagonal parameters only, which is shared with the augmented SDE.
@@
 The behavioral decoder uses two parallel linear transformations on the latent states to generate the mean and squared variance of a Gaussian distribution at each time step, defining the predicted behavior distribution.
+4 Baseline Model Specifications
+To ensure a fair comparison across all models, we fix the encoder and decoder architectures while varying only the latent sequential dynamics model.
+This section provides comprehensive architectural details and parameter counts for each model variant.
+Training Configuration
+All models (except gpSLDS) were trained using the hyperparameters specified in Table 2. Key training details include:
+All models trained via variational inference with ELBO objective
+Probabilistic initial state: where and are produced by the shared observation encoder
+For SDE models: Euler-Maruyama solver with step size
+For ODE models: Euler solver with same step size
+For RNN/GRU/LSTM: Standard discrete-time updates at observation time points
+Cyclic KL annealing applied to all variational models
+Implementation Notes
+Neural ODE & Neural SDE
+These models use flexible neural network parameterizations for the drift (and diffusion) functions.
+The MLP architecture allows the model to learn arbitrary smooth dynamics while maintaining computational tractability through modern ODE/SDE solvers.
+Coupled Oscillator Models (CO-ODE & CO-SDE)
+The coupled oscillator models incorporate mechanistic priors about neural oscillations. Each oscillator evolves according to:
+where the coupling strengths are modulated by the stimulus through a small neural network. This inductive bias significantly reduces the parameter count while maintaining expressiveness.
+Recurrent Models (RNN, GRU, LSTM)
+These discrete-time models are evaluated at observation time points. To ensure fair comparison with continuous-time models, we:
+Sample hidden states at the same time points as ODE/SDE models
+Use the same batch size and training duration
+Apply the same variational inference framework with probabilistic initial states
+Fair Comparison Justification
+By fixing encoder and decoder architectures, we isolate the impact of different latent dynamics models on performance.
+Any differences in results can be attributed solely to how well each model captures the temporal dynamics of neural and behavioral data.
+The parameter counts in Table 1 of the main text reflect only the latent dynamics parameters, as the encoder/decoder parameters (shared across all models) do not contribute to performance differences.
+Evaluation Metrics
+Bits per Spike
+We assess the quality of neural response predictions using the bits per spike (bps) metric, an information-theoretic quantity that measures the gain in predictive log-likelihood of a model over a time-homogeneous Poisson baseline, normalized by the number of spikes.
+The metric is grounded in the Poisson log-likelihood and has been widely adopted in recent neural latent variable modeling literature.
+Let denote the observed spike count for neuron at time bin , and let denote the firing rate predicted by the model.
+Let denote the empirical mean firing rate of neuron over the evaluation data.
+The bits per spike is computed as:
+This expression reflects the difference in log-likelihood (in bits) between the model and a baseline that predicts a constant firing rate for each neuron.
+Positive values indicate better predictive performance than the baseline; zero corresponds to parity; negative values indicate worse-than-baseline predictions.
+Our use of this metric is inspired by [1], who introduced bits per spike as a normalized log-likelihood score in the context of co-smoothing, wherein the goal is to predict the activity of held-out neurons conditioned on held-in neurons on the same trial.
+In contrast, we apply this metric to all recorded neurons on entirely leave-out trials that are not seen during training.
+This choice evaluates the model’s ability to generalize temporally, rather than across neurons, and aligns with our objective of modeling neural dynamics over continuous time.
+Coefficient of Determination ()
+To evaluate the model’s ability to predict behavioral measurements (e.g., kinematic variables or choice-related signals), we use the coefficient of determination, , under a Gaussian readout from the inferred latent state.
+This metric measures the proportion of variance in the observed behavior explained by the model’s predictions.
+Let be the observed behavioral measurement at time , and the corresponding model prediction. The score is defined as:
+where is the empirical mean of over the evaluation set.
+An score of 1 indicates perfect prediction, 0 corresponds to performance equal to predicting the mean, and negative values reflect worse-than-mean predictions.
+This metric provides an interpretable measure of behavioral predictability from latent neural dynamics.
+Supplementary References